### v3

# MindVote: When AI Meets the Wild West of Social Media Opinion

Abstract
Large Language Models (LLMs) are increasingly used as scalable tools for pilot testing, predicting public opinion distributions before deploying costly surveys. To serve as effective pilot testing tools, the performance of these LLMs is typically benchmarked against their ability to reproduce the outcomes of past structured surveys. This evaluation paradigm, however, is misaligned with the dynamic, context-rich social media environments where public opinion is increasingly formed and expressed. By design, surveys strip away the social, cultural, and temporal context that shapes public opinion, and LLM benchmarks built on this paradigm inherit these critical limitations. To bridge this gap, we introduce MindVote, the first benchmark for public opinion distribution prediction grounded in authentic social media discourse. MindVote is constructed from 3,918 naturalistic polls sourced from Reddit and Weibo, spanning 23 topics and enriched with detailed annotations for platform, topical, and temporal context. Using this benchmark, we conduct a comprehensive evaluation of 15 LLMs. MindVote provides a robust, ecologically valid framework to move beyond survey-based evaluations and advance the development of more socially intelligent AI systems.
Keywords Opinion Prediction  $\cdot$ Large Language Models  $\cdot$ Social Media  $\cdot$ Cross-Cultural Evaluation  $\cdot$ LLM Benchmark
1 Introduction
Figure 1: MindVote benchmark addresses three key limitations of previous survey-based approaches. Our benchmark provides diverse topics, cultural integration, and rich contextual metadata, overcoming the topic imbalance, cultural homogeneity, and context poverty of traditional survey datasets.
A core application for Large Language Models (LLMs) in understanding public opinion is serving as a rapid, scalable tool for pilot testing—predicting how a population will respond to a question before deploying costly, large-scale surveys rothschild2024opportunities ; bisbee2024synthetic ; suh2025language ; sinacola2025llms ; cao2025specializing ; anthis2025llm ; binz2025should . This capability is crucial across a wide spectrum of domains. For example, LLMs can be used to gauge reactions to new entertainment content and technological innovations, understand career aspirations, navigate complex social issues, and track evolving lifestyle trends argyle2023out ; qian2025large ; chen2025large . However, the central goal is not only forecasting the majority view, but also accurately predicting the entire public opinion distribution across a range of choices because the shape of the distribution reveals crucial information about societal consensus, polarization, and the intensity of dissent, which is lost in a single majority statistic zhou-etal-2022-distributed ; moon-etal-2024-virtual ; meister-etal-2025-benchmarking ; cao-etal-2025-specializing .
The prevailing paradigm for public opinion distribution prediction relies on traditional structured surveys krogstad2016pew ; santurkar2023whose ; elkjaer2024estimating ; suh2025language . This methodology is misaligned with the fast-paced, context-rich digital environments where opinions are now actively formed and expressed, such as social media platforms (boutyline2021cultural, ; riva2022social, ). While surveys include basic demographic questions—such as age, gender, or education level—these static variables are poor proxies for the dynamic, situated nature of social context zaller1992simple . Demographics may indicate who a person is in categorical terms, but they fail to reflect how opinions are shaped by lived experiences, group affiliations, ongoing conversations, or culturally specific reasoning patterns turner1979social ; kahan2011cultural ; markus2014culture ; mcintosh2019exploring .
Consequently, LLM benchmarks built on this flawed paradigm, such as OpinionQA and SubPop (santurkar2023whose, ; suh2025language, ), inherit these critical limitations. This reveals a dangerous gap in the responsible deployment of LLMs. As these models are increasingly used to act as a proxy for human populations, their accuracy becomes paramount. A proper evaluation framework must identify when models fail at context-rich, real-world tasks. Without this, we risk promoting systems that misrepresent the groups they are meant to simulate. This misrepresentation has severe downstream consequences, potentially compromising strategic decisions across critical domains like public policy, marketing, technology development, and community management (qi2024impact, ; radivojevic2024llms, ; qu2024performance, ; panickssery2024llm, ; repec:arx:papers:2502.03158, ; wang2025large, ). This reliance on traditional survey-based evaluation creates three critical gaps in our ability to assess an LLM’s true social understanding:
The Gap about Topic Imbalance. Real-world online discourse is thematically diverse tran-ostendorf-2016-characterizing ; song2023understanding ; entertainment, for instance, drives 70% of traffic on Weibo (chinamarketingcorp2024weibo, ), and Reddit’s largest communities focus on gaming and technology (proferes2021studying, ). Yet, current survey-based benchmarks are heavily skewed towards formal, institutional topics like politics, which represent only a fraction of naturalistic opinion expression (santurkar2023whose, ; zhao-etal-2024-worldvaluesbench, ). This mismatch means we cannot evaluate a model’s ability to adapt its reasoning from the formal discourse of a political survey to the vernacular social media discourse karjus2024evolving ; reveilhac2024augmenting .
The Gap about Cultural Homogeneity. Existing cross-cultural survey benchmarks often exhibit cultural homogeneity, creating test conditions on Western-centric questions even with other languages translation (haerpfer2022world, ; durmus2024towards, ; zhao-etal-2024-worldvaluesbench, ). A model might perform well on a translated question, but we are left unable to determine if this success stems from genuine cultural understanding or a superficial linguistic competence singh-etal-2025-global .
The Gap about Social Context Poverty. Opinions on social media are not noise but actual signals for prediction lopez2021framing ; they are primed by platform norms (e.g., Reddit’s anonymity vs. Weibo’s public-facing profiles (brown2018reddit, ; zhu2024privacy, )), temporal events banducci2015surveys , and community-specific discourse (tan2016winning, ; qiao2023does, ). Survey-based benchmarks systematically remove this contextual priming to achieve standardization (tourangeau2003context, ), preventing a true evaluation of a model’s ability to leverage these critical predictive cues 10.24963/ijcai.2023/749 .
To overcome these limitations, we extend the opinion distribution prediction scope of survey to a source of direct public opinion formation scarano2024analyzing : naturalistic social media polls. These polls directly address the three critical gaps, providing a robust foundation for testing the very abilities that current benchmarks obscure. The comparison between existing survey-based benchmarks and our benchmark is in Figure 1.
Figure 2: Distribution of topics in the MindVote dataset across five major topics and 23 subtopics. Percentages indicate the relative frequency of each subtopic.
Our Contributions. We introduce MindVote, the first benchmark for public opinion distribution prediction grounded in authentic social media discourse. We advance the field with the following innovations:
  • We construct and release MindVote, a benchmark of 3,918 authentic polls from two platforms (Reddit and Weibo) in their native English and Chinese. It spans 5 major topics and 23 sub-topics (shown in Figure 2 and is enriched with detailed social context annotations served as context priming.
  • We establish performance baselines for leading LLMs, identifying the best-performing models while also uncovering that models specifically fine-tuned on traditional surveys reveal a critical “survey-based specialization pitfall."
  • We provide strong evidence that enhancing a model’s capacity for in-context social reasoning is a more robust and effective path forward than fine-tuning on context-stripped data, offering a new direction for developing more socially intelligent systems.
2 Related Work
Opinion Distribution Prediction Benchmarks. U.S.-focused benchmarks like OpinionQA (santurkar2023whose, ) and SubPop (suh2025language, ) concentrate on political topics, while international efforts like GlobalOpinionQA (durmus2024towards, ) is Western-centric frameworks and WorldValuesBench (zhao-etal-2024-worldvaluesbench, ) is framed as multi-cultural value prediction. cao2025specializing introduces three datasets (two English, one Chinese) specifically for group-level distribution prediction. These approaches contain only demographic-value pairings, structured survey format, or Western-centric lens that abstract away the naturalistic cultural contexts and context priming that are essential for authentic opinion distribution prediction.
Figure 3: This flow pipeline demonstrates the dataset construction from data source to final composition of datasets.
3 Benchmarking Setup
3.1 Dataset Construction
MindVote’s construction involved transforming 7,648 raw polls into 3,918 high-quality polls through strategic data sourcing, social context annotation, rigorous quality control, and the cross-cultural augmentation. Figure 3 shows the flow of dataset construction.
Platform Selection Strategy. Our platform selection strategy is designed to evaluate distinct aspects of LLM opinion prediction capabilities. Reddit provides an anonymous environment where users express unfiltered, authentic opinions without identity-based social constraints, requiring models to interpret and predict genuine thoughts purely through content analysis without relying on user profiles or reputation cues. Weibo enables the evaluation of model performance in culture-specific contexts, testing models’ understanding of Chinese cultural contexts, social norms, and culture-laden discourse patterns.
Multi-Platform Data Collection. We collected 7,648 polls across three platforms spanning 2019-2025. Weibo dataset contributed 3,757 polls (2,026 from existing datasets lu-etal-2021-engage from 2019 to 2021, 1,734 newly crawled) and are anonymized for users’ personal identifiable information, capturing Chinese social media dynamics across pandemic and post-pandemic periods. Reddit provided 3,891 polls from diverse subreddits including r/poll during 2021-2025.
Quality Control Pipeline. Our four-stage pipeline efficiently produced high-quality 1,959 core polls:
  1. 1.
    Initial Preprocessing: This process included the removal of commercial votes, targeting promotional polling content that lacked authentic user engagement. Furthermore, format-corrupted poll data was identified and discarded through a systematic validation of structural integrity.
  2. 2.
    Content Quality Filtering: This filtering begins with duplicate content removal using a MinHash algorithm to eliminate items with greater than 95% overlap (shrivastava2014defense, ), effectively targeting redundant trending topics, reposted polls, and cross-platform duplicates. Subsequently, automated toxicity screening was conducted using the Google Perspective API, with polls retaining a toxicity score below 0.4 being retained to filter out hate speech and overly controversial subjects (lees2022new, ).
  3. 3.
    Social Relevance Validation: Human verification of voting patterns (removing polls $\leq$ 100 votes for social relevance) and verification of poll and vote existence in all three platforms to ensure authentic social engagement and meaningful community participation.
  4. 4.
[…]
We manually annotate each poll includes rich metadata for social context enrichment: general platform context (e.g., user statistics and user behaviors), topical context within that platform (topic-specific discourse patterns), and temporal context for each poll (social events and backgrounds related to the poll when created).
Cross-Cultural Augmentation. We created a parallel bilingual corpus for all three platforms through machine translation with rigorous quality control to demonstrate both linguistic and cultural effects: 10% back-translation validation (BLEU $>$ 35 (papineni2002bleu, )), 5% native speaker rewriting, and expert review (Fleiss’ $\kappa$ = 0.51). The total number of polls becomes 3,918 after augmentation.
Final Dataset Composition. The final MindVote dataset is composed of 3,918 polls, with 2,158 sourced from Reddit and 1,760 from Weibo. Each poll is enriched with a comprehensive set of metadata, including its creation time, total vote count, and three layers of social context: general platform context, topical context within that platform, and the specific temporal context at the time the poll was created. To ensure broad accessibility and ease of use, the entire dataset is provided in both CSV and JSON formats. Table 1 shows one example with its simplified metadata keywords for illustration.
Table 1: An example poll from our MindVote dataset, demonstrating the structure of the question, options, and associated social context provided for model evaluation.
Poll: How threatened do you feel by AI replacing your job?
(Platform: Reddit, Date: April 14, 2025, Votes: 6,567)
Options 1. Not at all threatened 2. Slightly threatened 3. Moderately threatened 4. Very threatened
Platform Context Reddit user base: 58% US, 46% college-educated, generally tech-oriented.
Topical Context Broad AI adoption. 78% of organizations use AI; 55% of Americans use AI regularly.
Temporal Context AI-driven layoff spike fuels job insecurity.
3.2 Experiment Design
Task. We evaluate LLMs on opinion distribution prediction within naturalistic social discourse in MindVote dataset including 3,918 polls. Given a poll question with metadata including social contexts, models predict the probability distribution over answer choices.
Models. We evaluate 15 leading LLMs across closed-source, open-source and specalization categories: Claude-3.7-Sonnet-02-24 (claude-3.7-thinking, ), GPT-4o-2024-11-20 (openai-4o, ), GPT-4.1-04-14 (openai-4.1, ), Gemini-2.5-Pro-05-06 (google-gemini-2-5-pro, ), o3-medium-04-16 (openai-o3-o4mini-system-card, ), Deepseek-R1-05-28 (2501.12948, ), Qwen2.5-32B-Instruct (qwen2.5, ), Gemma-2-9B-it (team2024gemma, ), Mistral-7B-Instruct-v0.1 chaplot2023albert , Llama-2-13B-Chat touvron2023llama , Llama-3-70B-Instruct (grattafiori2024llama, ), and Llama-4-Maverick-17B-128E (llama4, ). We also evaluate three opinion distribution prediction specialization models from suh2025language . These models are fine-tuned using LoRA hu2022lora on base models: Llama-3-70B-Base grattafiori2024llama , Llama-2-13B-Base touvron2023llama , and Mistral-7B-v0.1 chaplot2023albert which we named them respectively as: SubPop-Llama-3-70B, SubPop-Llama-2-13B, SubPop-Mistral-7B.
Pipeline. All primary evaluations use a greedy decoding strategy (temperature=0) with default hyperparameter settings under zero-shot with context priming. For the specialization fine-tuned models, we adapt those models into our pipeline by first loading the respective pretrained base model and then applying the publicly available LoRA weight checkpoints provided by suh2025language .
Prompt. To ensure consistent and machine-readable outputs, we employ a structured prompting strategy where the model is given a JSON object containing the poll and its context. The template instructs the model to assume the role of a “opinion distribution prediction expert analyzing voting patterns and social dynamics." The prompt includes the poll question and is enriched with social context metadata served as context priming, with instruction for step-by-step reasoning. The model’s task is to return a JSON object with a schema identical to the input, replacing placeholder fields with its numeric predictions for the voting distribution. The complete and structured template is detailed in Appendix LABEL:appendix:prompt.
Evaluation Metrics. We adopt four distinct metrics to provide a comprehensive evaluation. Our primary metric is 1 - Wasserstein Distance (1-Wass.) (santurkar2023whose, ; meister-etal-2025-benchmarking, ; suh2025language, ). The Wasserstein Distance measures the minimum cost for transforming one distribution into another, crucially accounting for semantic similarity between answer choices by treating them as points in a metric space. To complement this, we also report Spearman’s Rank Correlation Coefficient ($\rho$) (zhou-etal-2022-distributed, ; pavlopoulos2023distance, ), a non-parametric measure of how well the predicted ranking of options matches the true ranking of vote shares; 1 - KL Divergence (meister-etal-2025-benchmarking, ; nguyen2025task, ), which quantifies the information loss when using the model’s predicted distribution to approximate the ground truth; and One-hot Accuracy (zhou-etal-2022-distributed, ; santurkar2023whose, ; suh2025language, ), which provides a strict measure of whether the single most likely predicted answer is correct.
Evaluation Boundary. We include upper bounds and lower bounds for comparisons following suh2025language . The upper bound is established by sampling subsets of the original results, calculating the four metrics between subsampled and original distributions, and performing bootstrapping to obtain a robust estimate that captures the intrinsic variance arising from the respondent sampling process in opinion. The uniform distribution lower bound establishes a performance floor equivalent to random chance.
Table 2: Opinion distribution prediction performance of LLMs on the MindVote Benchmark. Scores are presented as Mean values, evaluated on four different metrics: 1-Wasserstein distance (1-Wass.), 1-KL Divergence (1-KL Div.), Spearman’s Rank Correlation (Spearman.), and One-hot Accuracy (Acc.). All metrics are the higher the better.
Model 1-Wass. 1-KL Div. Spearman. Acc.
Closed-source Models
o3-medium 0.892 0.859 0.756 0.581
Gemini-2.5-Pro 0.891 0.845 0.751 0.564
Claude-3.7-Sonnet 0.891 0.851 0.722 0.551
GPT-4o 0.880 0.836 0.691 0.711
GPT-4.1 0.874 0.845 0.688 0.697
Open-source Models
Deepseek-R1 0.876 0.831 0.739 0.558
Qwen2.5-32B 0.866 0.787 0.605 0.483
Llama-4-17B 0.820 0.731 0.659 0.429
Llama-3-70B 0.844 0.752 0.641 0.461
Llama-2-13B 0.807 0.718 0.592 0.369
Gemma-2-9B 0.802 0.705 0.575 0.362
Mistral-7B 0.808 0.719 0.597 0.365
Specialization Models
SubPop-Llama-3-70B 0.805 0.713 0.593 0.417
SubPop-Llama-2-13B 0.774 0.693 0.558 0.378
SubPop-Mistral-7B 0.782 0.695 0.546 0.370
Upper Bound 0.972 0.976 0.961 0.964
Lower Bound 0.701 0.663 0.000 0.307
4 Result
4.1 Performance of General Purpose Models
We analyze the overall performance of ten leading general purpose LLMs on the MindVote benchmark in Table 2. Across all four metrics, the closed-source model o3-medium consistently outperforms all other models, establishing the highest performance ceiling among current general-purpose LLMs. Among open-source models, Deepseek-R1 demonstrates the strongest results, narrowly outperforming other models in its class. However, a critical gap remains between the top-performing models and the upper bound for all metrics. This highlights substantial opportunities for improvement in modeling authentic public opinion distribution.
4.2 Survey-based Specialization Pitfalls
Our result in Table 2 reveals that fine-tuning on a structured survey dataset results in a significant performance drop on authentic, in-domain tasks. To demonstrate this, we compare our specialization models (base models that fine-tuned on traditional but context-stripped survey data) against their original base instruction-tuned models (the general-purpose models).
As shown in Table 2, the specialization models consistently underperform their base counterparts. This evaluation was conducted under our standard testing condition, where all models were provided with full context served as context priming setting. For instance, the SubPop-Llama-2-13B specialization model exhibits a 3.3 percentage point drop in its 1-Wass. score compared to the base Llama-2-13B model. This finding suggests that fine-tuning on sanitized data fails to equip models with the necessary capabilities to handle real-world contextual complexity. Instead of improving performance, this narrow specialization is detrimental, degrading the models’ ability to generalize to authentic social discourse.
4.3 Performance Across Topics and Cultures
Performance varies by Topic. As shown in Table 3, model performance is not uniform across different topical domains. On average, models achieve higher performance on Social Issues and Lifestyle topics, which often feature more structured discourse patterns that align with standard pre-training corpora. In contrast, performance tends to be lower in domains like Technology and Entertainment where specialized jargon, community-specific norms, and rapidly evolving vernacular are more common. This disparity suggests that models struggle to adapt their reasoning from general knowledge to the specific linguistic and social conventions of diverse online communities.
Performance Reflects Cultural Origin. The analysis in Figure 4 reveals a ’Cultural Gap,’ showing models have a strong cultural alignment with their origin. Western models like Gemini-2.5-Pro excel on Reddit, while Chinese models such as DeepSeek-R1 dominate on Weibo. We confirmed this gap is primarily cultural, not linguistic, by evaluating translated content; the performance penalty from translation was minimal compared to the large drop across cultural sources. This systematic variation demonstrates a clear alignment effect, where models are more adept at interpreting the cultural norms of their training data.
4.4 Analysis
Section 4.2 reveals a critical survey-based specialization pitfall: base models fine-tuned on sanitized survey datasets exhibit degraded performance even with context priming. This unexpected finding raises three questions that we expect to investigate:
  • (1) What role does social context actually play?
  • (2) Why do models struggle with contextual complexity?
  • (3) Is it more beneficial for a model to understand specific context or to imitate examples?
These questions emerge naturally from our specialization findings. The failure of fine-tuned models despite having full context suggests either that platform, topic and temporal context are less important than assumed, or that these models cannot effectively utilize those contextual information. This motivates our first investigation into whether social context is truly beneficial signal or noise. The stark performance gap between base and specialized models further suggests that certain contextual characteristics may be particularly challenging, leading to our second analysis of contextual complexity dimensions. Finally, the contrast between our context-rich approach and traditional few-shot learning methods—where specialized models were trained on examples rather than contextual understanding—compels us to directly compare these two information-providing strategies.
Table 3: Opinion distribution prediction performance across different topics. The final row shows the average performance across all models for each major topic.
Model Tech. Soc. Enter. Career Life.
Closed-source Models
o3-medium 0.868 0.925 0.869 0.904 0.903
Gemini-2.5-Pro 0.860 0.917 0.881 0.896 0.906
Claude-3.7-S. 0.865 0.908 0.896 0.891 0.894
GPT-4o 0.859 0.893 0.872 0.895 0.884
GPT-4.1 0.866 0.913 0.854 0.878 0.868
Open-source Models
Deepseek-R1 0.889 0.892 0.868 0.879 0.887
Qwen2.5-32B 0.846 0.876 0.867 0.868 0.875
Llama-4-17B 0.824 0.844 0.832 0.866 0.857
Llama-3-70B 0.835 0.853 0.836 0.845 0.854
Llama-2-13B 0.798 0.824 0.801 0.789 0.817
Gemma-2-9B 0.802 0.839 0.832 0.848 0.885
Mistral-7B 0.809 0.836 0.812 0.804 0.828
Specialization Models
SubPop-Llama-3-70B 0.794 0.821 0.798 0.786 0.813
SubPop-Llama-2-13B 0.762 0.798 0.769 0.751 0.785
SubPop-Mistral-7B 0.771 0.803 0.778 0.764 0.792
Average 0.830 0.863 0.838 0.844 0.857
Social Context is a Critical Signal. Our context ablation study (Table 4) confirms that social context is not noise but a vital signal for accurate opinion prediction. Removing all contextual metadata (No Context) results in the most substantial performance drop on average (-5.91%), followed closely by platform-specific information removal (-5.12%) and topical context removal (-4.52%). Temporal context shows the smallest impact, but its existence is important to provide backgrounds for questions need temporal information.
Notably, specialized fine-tuned models exhibit dramatically larger degradation across all ablation conditions. This severe degradation approaches the theoretical lower bound of uniform distribution performance, suggesting these models have become overly dependent on contextual signals. Since these models were fine-tuned on structured survey data in standardized formats, they struggle to adapt when deployed on natural social media polls when contextual information may be fragmented or entirely absent. While these models may excel in controlled survey environments, their heightened sensitivity to context removal in real-world social media settings reveals a limitation, potentially constraining their applicability across diverse online environments.
Figure 4: LLMs exhibit a cultural alignment with their origin. Models trained from Western data better on Reddit (EN-English), while those from China (e.g., DeepSeek-R1) excel on Weibo (ZH-Chinese). This cultural gap persists after controlling for linguistic effects. Error bars represent 95% CI.
Contextual Complexity. Our ablation study showed that removing social context severely degrades performance. To understand why this context is so critical, we investigate how its inherent complexity impacts model behavior. We define contextual complexity using three metrics: (1) character length of combined metadata; (2) language informality (measured by slang and emoji frequency); and (3) topic niche nature (inversely proportional to dataset percentage). As demonstrated in Figure 5, our findings reveal consistent negative correlations between model performance and all complexity metrics across evaluated systems. This indicates that as social contexts become longer, more informal, and more specialized, models systematically exhibit reduced prediction certainty. The analysis reveals distinct performance tiers, with closed-source models showing greater resilience, open-source models displaying moderate sensitivity, and specialization models exhibiting the most pronounced degradation under complexity. This pattern suggests that domain-specific fine-tuning on formal survey data creates brittleness when encountering the informal, less structured nature of social media polls. The hierarchical performance differences confirm that authentic social discourse’s rich, messy characteristics pose universal challenges for LLMs, with models fine-tuned on structured formats being particularly vulnerable to contextual complexity.
Table 4: Opinion distiribution prediction performance degradation from the full-context baseline. All scores represent the drop in 1-Wasserstein Distance (%). Abbreviations: Plat. (Platform), Topic. (topical), Temp. (Temporal), Ctx. (Context).
Model w/o Plat. w/o Topic. w/o Temp. No Ctx.
Closed-source Models
o3-medium -3.03 -2.31 -1.02 -4.13
Gemini-2.5-Pro -2.37 -3.30 -0.98 -4.34
Claude-3.7-Sonnet -4.46 -3.91 -1.39 -5.53
GPT-4o -5.75 -3.39 -0.87 -4.92
GPT-4.1 -4.92 -3.36 -1.57 -5.30
Open-source Models
Deepseek-R1 -4.19 -3.55 -1.04 -5.98
Qwen2.5-32B -4.42 -3.81 -0.88 -6.24
Llama-3-70B -4.82 -4.78 -2.11 -5.19
Llama-4-17B -5.80 -5.33 -1.01 -6.08
Llama-2-13B -4.95 -5.26 -1.37 -5.92
Gemma-2-9B -5.37 -4.74 -1.26 -6.47
Mistral-7B -5.92 -4.84 -1.46 -6.55
Specialization Models
SubPop-Llama-3-70B -6.82 -6.68 -1.88 -6.95
SubPop-Llama-2-13B -6.94 -6.11 -2.10 -7.23
SubPop-Mistral-7B -6.89 -6.41 -2.96 -7.54
Average -5.12 -4.52 -1.46 -5.91
Contextual Priming vs. Few-Shot Learning. Our “survey-based specialization pitfall" finding (Section 4.2) raised a key question: do models benefit more from understanding a poll’s environment via contextual priming or from imitating examples via few-shot learning? To test this, we compared model performance in a zero-shot setting with full context against a zero-shot setting with no context and a few-shot setting with no context.
The results in Figure 6 are stark: contextual priming yields a substantial increase across all models, exceeding the impact of few-shot examples in most cases. Moreover, few-shot performance shows instability across models but contextual priming provides more consistent improvements. This finding suggests that the ability to situate a problem within its social context is a more critical capability than pattern-matching from isolated examples.
Figure 5: Performance degradation patterns across model types and complexity dimensions. The heatmap reveals a clear hierarchy: closed-source models maintain robustness while survey-specialized models exhibit severe brittleness to social media contextual complexity.
5 Error Analysis
We conduct an error analysis to categorize prediction reasoning errors  [^1] We consider the individual poll’s 1-Wass. $<$ 0.8 as error.. Using the LLM-as-a-judge, we identify three main categories of error type shown in Table 5.
Figure 6: Contextual priming outperforms few-shot learning across all model categories, highlighting the importance of social context over example-based pattern matching.
Table 5: Analysis of Claude 3.7 Sonnet prediction errors categorized by failure modes: Platform Mis-Adaptation (42.5%) – misapplying knowledge about specific platforms; Cultural Misalignment (36.6%) – ignoring local and cultural contexts; Temporal Dislocation (20.0%) – mis-understanding about the temporal context.
Error Type Case Study & Failed Reasoning
Platform Misadaptation Influencer Subscription Worth or Not (Reddit): Failing to recognize Reddit attitude toward ad-centric monetization and user engagement patterns.
Cultural Misalignment New phone Buy or Wait (Weibo): Assumed universal tech-enthusiasm drives upgrades, overlooking Chinese market’s specific socio-economic factors and consumer behavior patterns.
Temporal Dislocation Calling it "X" vs. "Twitter" (Reddit): Reasoning anchored to official rebrand timeline, underestimating persistent colloquial usage and real-world adoption resistance.
Platform Mis-Adaptation. A model can fail into platform mis-adaptation, where they misapplying their knowledge about certain platform, such as when analyzing a Reddit poll about paid TikTok subscriptions. It may apply a generic economic lens, predicting high tendency toward worth option based on the influencer economy. This reasoning ignores the context of Reddit, where the user tend to be negative of influencer monetization. The prediction is thus inverted from the ground truth which shows a much higher proportion of not worth it.
Cultural Misunderstanding. In the phone replacement case, models systematically misinterpret Chinese users’ pragmatic spending attitudes, failing to capture the genuine cost-conscious preferences common among Weibo users, which often contradict Western-centric consumer marketing narratives. While model recognizes about the technical enthusiasm in Weibo, their predictions fail to reflect the cost-conscious approach that characterizes technology adoption for many Chinese social media users.
Temporal Dislocation. This error happens when models fail to account for the lag between official changes and actual user adoption patterns. In the Twitter-to-X rebranding case, the model assumes that users would quickly adopt the new term. However, this prediction overlooks that Reddit users, in particular, demonstrated strong attachment to the familiar terminology, viewing continued use of the old name as both habitual behavior and subtle resistance. The model’s temporal reasoning thus misaligned with the gradual and reluctant pace of real-world language adoption.
6 Conclusion
We introduce MindVote, the first benchmark for public opinion prediction grounded in authentic social media discourse. Our comprehensive evaluation demonstrates the critical importance of assessing models in naturalistic, context-rich environments. We argue that the path to socially intelligent AI requires enhancing a model’s capacity for in-context reasoning. Our results demonstrate that models perform best when they can explicitly identify, weigh, and interpret the social cues present in the immediate context—a skill that requires flexible reasoning rather than memorized associations. MindVote provides the essential tool to guide and measure this necessary shift.

### v5

@@
 Large Language Models (LLMs) are increasingly used as scalable tools for pilot testing, predicting public opinion distributions before deploying costly surveys.
-To serve as effective pilot testing tools, the performance of these LLMs is typically benchmarked against their ability to reproduce the outcomes of past structured surveys.
-This evaluation paradigm, however, is misaligned with the dynamic, context-rich social media environments where public opinion is increasingly formed and expressed.
-By design, surveys strip away the social, cultural, and temporal context that shapes public opinion, and LLM benchmarks built on this paradigm inherit these critical limitations.
-To bridge this gap, we introduce MindVote, the first benchmark for public opinion distribution prediction grounded in authentic social media discourse.
-MindVote is constructed from 3,918 naturalistic polls sourced from Reddit and Weibo, spanning 23 topics and enriched with detailed annotations for platform, topical, and temporal context.
-Using this benchmark, we conduct a comprehensive evaluation of 15 LLMs.
+However, the prevailing paradigm for evaluating these models relies on traditional structured surveys—a methodology misaligned with the more realistic scenarios like social media where opinions are rich in digital contexts.
+By design, surveys strip away the social and cultural context that shapes public opinion, and LLM benchmarks built on this paradigm inherit these critical limitations.
+To bridge this gap, we introduce MindVote, the first benchmark for public opinion prediction grounded in authentic social media discourse.
+MindVote is constructed from 3,918 naturalistic polls sourced from Reddit and Weibo, spanning 23 topics and enriched with detailed annotations for platform and topical context.
+Using this benchmark, we conduct a comprehensive evaluation of 15 LLMs, revealing a critical “survey-based specialization pitfall” where models fine-tuned on traditional surveys underperform their general-purpose counterparts and demonstrating the necessity of context in social media.
 MindVote provides a robust, ecologically valid framework to move beyond survey-based evaluations and advance the development of more socially intelligent AI systems.
-Keywords Opinion Prediction  $\cdot$ Large Language Models  $\cdot$ Social Media  $\cdot$ Cross-Cultural Evaluation  $\cdot$ LLM Benchmark
 1 Introduction
@@
 Our benchmark provides diverse topics, cultural integration, and rich contextual metadata, overcoming the topic imbalance, cultural homogeneity, and context poverty of traditional survey datasets.
-A core application for Large Language Models (LLMs) in understanding public opinion is serving as a rapid, scalable tool for pilot testing—predicting how a population will respond to a question before deploying costly, large-scale surveys rothschild2024opportunities ; bisbee2024synthetic ; suh2025language ; sinacola2025llms ; cao2025specializing ; anthis2025llm ; binz2025should .
+A core application for Large Language Models (LLMs) in understanding public opinion is serving as a rapid, scalable tool for pilot testing—predicting how a population will respond to a question before deploying costly, large-scale surveys (Rothschild et al.
+2024; Bisbee et al.
+2024; Suh et al.
+2025; Sinacola, Pachot, and Petit 2025; Cao et al.
+2025a; Anthis et al.
+2025; Binz et al.
+2025).
 This capability is crucial across a wide spectrum of domains.
-For example, LLMs can be used to gauge reactions to new entertainment content and technological innovations, understand career aspirations, navigate complex social issues, and track evolving lifestyle trends argyle2023out ; qian2025large ; chen2025large .
-However, the central goal is not only forecasting the majority view, but also accurately predicting the entire public opinion distribution across a range of choices because the shape of the distribution reveals crucial information about societal consensus, polarization, and the intensity of dissent, which is lost in a single majority statistic zhou-etal-2022-distributed ; moon-etal-2024-virtual ; meister-etal-2025-benchmarking ; cao-etal-2025-specializing .
-The prevailing paradigm for public opinion distribution prediction relies on traditional structured surveys krogstad2016pew ; santurkar2023whose ; elkjaer2024estimating ; suh2025language .
-This methodology is misaligned with the fast-paced, context-rich digital environments where opinions are now actively formed and expressed, such as social media platforms (boutyline2021cultural, ; riva2022social, ).
-While surveys include basic demographic questions—such as age, gender, or education level—these static variables are poor proxies for the dynamic, situated nature of social context zaller1992simple .
-Demographics may indicate who a person is in categorical terms, but they fail to reflect how opinions are shaped by lived experiences, group affiliations, ongoing conversations, or culturally specific reasoning patterns turner1979social ; kahan2011cultural ; markus2014culture ; mcintosh2019exploring .
-Consequently, LLM benchmarks built on this flawed paradigm, such as OpinionQA and SubPop (santurkar2023whose, ; suh2025language, ), inherit these critical limitations.
-This reveals a dangerous gap in the responsible deployment of LLMs.
-As these models are increasingly used to act as a proxy for human populations, their accuracy becomes paramount.
+For example, LLMs can be used to predict reactions to new entertainment content and technological innovations, understand career aspirations, navigate complex social issues, and track evolving lifestyle trends (Argyle et al.
+2023; Qian et al.
+2025; Chen, Wang, and Ma 2025).
+However, the central goal is not simply to identify the majority opinion, but to predict how opinions are distributed across all possible choices.
+Understanding the full distribution is important because it reveals whether society is unified, divided, or contains strong dissent—information that a single majority number cannot capture.
+(Zhou, Nie, and Bansal 2022; Moon et al.
+2024; Meister, Guestrin, and Hashimoto 2025; Cao et al.
+2025b).
+However, the prevailing paradigm for public opinion distribution prediction relies on traditional structured surveys (Krogstad, Passel, and Cohn 2016; Santurkar et al.
+2023; Elkjær and Wlezien 2024; Suh et al.
+2025).
+This approach does not align with the fast-paced, context-rich digital environments—such as social media platforms—where opinions are now actively formed and expressed (Boutyline and Soter 2021; Riva, Aureli, and Silvestrini 2022).
+While surveys include basic demographic questions—such as age, gender, or education level—these static variables are poor proxies for the dynamic, situated nature of social context (Zaller and Feldman 1992).
+Demographics may indicate who a person is in broad categories, but fail to reflect how opinions are specifically shaped by lived experiences, group affiliations, ongoing conversations, or culturally reasoning patterns (Turner, Brown, and Tajfel 1979; Kahan, Jenkins-Smith, and Braman 2011; Markus and Kitayama 2014; McIntosh and Wright 2019).
+Consequently, LLM benchmarks built on this flawed paradigm, such as OpinionQA and SubPop (Santurkar et al.
+2023; Suh et al.
+2025), inherit these critical limitations.
+This reveals a crucial gap in the responsible deployment of LLMs.
+As these models are increasingly used to act as a proxy for human populations, ensuring their accuracy becomes essential.
 A proper evaluation framework must identify when models fail at context-rich, real-world tasks.
 Without this, we risk promoting systems that misrepresent the groups they are meant to simulate.
-This misrepresentation has severe downstream consequences, potentially compromising strategic decisions across critical domains like public policy, marketing, technology development, and community management (qi2024impact, ; radivojevic2024llms, ; qu2024performance, ; panickssery2024llm, ; repec:arx:papers:2502.03158, ; wang2025large, ).
+This misrepresentation has severe downstream consequences, potentially compromising strategic decisions across critical topics like public policy, marketing, technology development, and community management (Qi 2024; Radivojevic, Clark, and Brenner 2024; Qu and Wang 2024; Panickssery, Bowman, and Feng 2024; Alekseenko et al.
+2025; Wang, Morgenstern, and Dickerson 2025).
 This reliance on traditional survey-based evaluation creates three critical gaps in our ability to assess an LLM’s true social understanding:
 The Gap about Topic Imbalance.
-Real-world online discourse is thematically diverse tran-ostendorf-2016-characterizing ; song2023understanding ; entertainment, for instance, drives 70% of traffic on Weibo (chinamarketingcorp2024weibo, ), and Reddit’s largest communities focus on gaming and technology (proferes2021studying, ).
-Yet, current survey-based benchmarks are heavily skewed towards formal, institutional topics like politics, which represent only a fraction of naturalistic opinion expression (santurkar2023whose, ; zhao-etal-2024-worldvaluesbench, ).
-This mismatch means we cannot evaluate a model’s ability to adapt its reasoning from the formal discourse of a political survey to the vernacular social media discourse karjus2024evolving ; reveilhac2024augmenting .
+Real-world online discourse is thematically diverse (Tran and Ostendorf 2016; Song et al.
+2023); for instance, entertainment drives 70% of traffic on Weibo (China Marketing Corp 2024), while Reddit’s largest communities focus on gaming and technology (Proferes et al.
+2021).
+Yet, current survey-based benchmarks are heavily skewed towards formal, institutional topics like politics, which represent only a fraction of naturalistic opinion expression (Santurkar et al.
+2023; Zhao et al.
+2024).
+Because of this mismatch, we cannot evaluate if a model can adjust its reasoning from formal political surveys to the everyday language used on social media (Karjus and Cuskley 2024; Reveilhac and Morselli 2024).
 The Gap about Cultural Homogeneity.
-Existing cross-cultural survey benchmarks often exhibit cultural homogeneity, creating test conditions on Western-centric questions even with other languages translation (haerpfer2022world, ; durmus2024towards, ; zhao-etal-2024-worldvaluesbench, ).
-A model might perform well on a translated question, but we are left unable to determine if this success stems from genuine cultural understanding or a superficial linguistic competence singh-etal-2025-global .
-The Gap about Social Context Poverty.
-Opinions on social media are not noise but actual signals for prediction lopez2021framing ; they are primed by platform norms (e.g., Reddit’s anonymity vs.
-Weibo’s public-facing profiles (brown2018reddit, ; zhu2024privacy, )), temporal events banducci2015surveys , and community-specific discourse (tan2016winning, ; qiao2023does, ).
-Survey-based benchmarks systematically remove this contextual priming to achieve standardization (tourangeau2003context, ), preventing a true evaluation of a model’s ability to leverage these critical predictive cues 10.24963/ijcai.2023/749 .
-To overcome these limitations, we extend the opinion distribution prediction scope of survey to a source of direct public opinion formation scarano2024analyzing : naturalistic social media polls.
-These polls directly address the three critical gaps, providing a robust foundation for testing the very abilities that current benchmarks obscure.
+Existing cross-cultural survey benchmarks often exhibit cultural homogeneity, creating test conditions on Western-centric questions even with other languages translation (Haerpfer et al.
+2022; DURMUS et al.
+2024; Zhao et al.
+2024).
+A model might perform well on a translated question, but we are left unable to determine if this success stems from genuine cultural understanding or a superficial linguistic competence (Singh et al.
+2025).
+The Gap about Context Poverty.
+Opinions on social media are not noise but are helpful predictive signals (López-Rabadán 2021).
+This is correlated with context priming, a phenomenon where exposure to a specific context influences how individuals perceive and respond (Doyle and Lee 2016).
+On social platforms, users are primed by factors such as platform norms (e.g., Reddit’s anonymity vs.
+Weibo’s public-facing profiles (Brown et al.
+2018; Zhu 2024)), temporal events (Banducci and Stevens 2015), and community-specific discourse (Tan et al.
+2016; Qiao 2023).
+In contrast, survey-based benchmarks systematically remove this contextual priming to achieve standardization (Tourangeau, Singer, and Presser 2003), which prevents a true evaluation of a model’s ability to leverage these critical predictive cues (Li et al.
+2023).
+To overcome these limitations, we move beyond surveys and instead use naturalistic social media polls—a more direct source for analyzing public opinion formation (Scarano et al.
+2024).
+These polls address three critical gaps and provide a robust foundation for evaluation.
 The comparison between existing survey-based benchmarks and our benchmark is in Figure 1.
@@
 Our Contributions.
-We introduce MindVote, the first benchmark for public opinion distribution prediction grounded in authentic social media discourse.
+We introduce MindVote, the first benchmark for public opinion distribution prediction grounded in realistic social media discourse.
 We advance the field with the following innovations:
   • We construct and release MindVote, a benchmark of 3,918 authentic polls from two platforms (Reddit and Weibo) in their native English and Chinese.
-It spans 5 major topics and 23 sub-topics (shown in Figure 2 and is enriched with detailed social context annotations served as context priming.
-  • We establish performance baselines for leading LLMs, identifying the best-performing models while also uncovering that models specifically fine-tuned on traditional surveys reveal a critical “survey-based specialization pitfall."
-  • We provide strong evidence that enhancing a model’s capacity for in-context social reasoning is a more robust and effective path forward than fine-tuning on context-stripped data, offering a new direction for developing more socially intelligent systems.
+It spans 5 major topics and 23 sub-topics (shown in Figure 2) and is enriched with detailed social context annotations.
+  • We benchmark leading LLMs, identify top performers, and reveal that models fine-tuned on traditional surveys face a critical survey-based specialization pitfall.
+  • We demonstrate that enhancing a model’s capacity with social-context reasoning is more effective than fine-tuning on context-stripped data, offering a new direction for developing more socially intelligent systems.
 2 Related Work
 Opinion Distribution Prediction Benchmarks.
-U.S.-focused benchmarks like OpinionQA (santurkar2023whose, ) and SubPop (suh2025language, ) concentrate on political topics, while international efforts like GlobalOpinionQA (durmus2024towards, ) is Western-centric frameworks and WorldValuesBench (zhao-etal-2024-worldvaluesbench, ) is framed as multi-cultural value prediction.
-cao2025specializing introduces three datasets (two English, one Chinese) specifically for group-level distribution prediction.
-These approaches contain only demographic-value pairings, structured survey format, or Western-centric lens that abstract away the naturalistic cultural contexts and context priming that are essential for authentic opinion distribution prediction.
+U.S.-focused benchmarks like OpinionQA (Santurkar et al.
+2023) and SubPop (Suh et al.
+2025) concentrate on political topics, while international efforts like GlobalOpinionQA (DURMUS et al.
+2024) is Western-centric frameworks and WorldValuesBench (Zhao et al.
+2024) is framed as multi-cultural value prediction.
+Cao et al.
+(2025a) introduces three datasets (two English, one Chinese) specifically for group-level distribution prediction.
+These approaches contain only demographic-value pairings, structured survey format, or Western-centric lens that abstract away the naturalistic cultural contexts that are essential for authentic opinion distribution prediction.
 Figure 3: This flow pipeline demonstrates the dataset construction from data source to final composition of datasets.
@@
 Multi-Platform Data Collection.
-We collected 7,648 polls across three platforms spanning 2019-2025.
-Weibo dataset contributed 3,757 polls (2,026 from existing datasets lu-etal-2021-engage from 2019 to 2021, 1,734 newly crawled) and are anonymized for users’ personal identifiable information, capturing Chinese social media dynamics across pandemic and post-pandemic periods.
+We collected 7,648 polls across two platforms spanning 2019-2025.
+Weibo dataset contributed 3,757 polls (2,026 from existing datasets (Lu et al.
+2021) from 2019 to 2021, 1,734 newly crawled) and are anonymized for users’ personal identifiable information, capturing Chinese social media dynamics across pandemic and post-pandemic periods.
 Reddit provided 3,891 polls from diverse subreddits including r/poll during 2021-2025.
@@
   2. 2.
-    Content Quality Filtering: This filtering begins with duplicate content removal using a MinHash algorithm to eliminate items with greater than 95% overlap (shrivastava2014defense, ), effectively targeting redundant trending topics, reposted polls, and cross-platform duplicates.
-Subsequently, automated toxicity screening was conducted using the Google Perspective API, with polls retaining a toxicity score below 0.4 being retained to filter out hate speech and overly controversial subjects (lees2022new, ).
+    Content Quality Filtering: This filtering begins with duplicate content removal using a MinHash algorithm to eliminate items with greater than 95% overlap (Shrivastava and Li 2014), effectively targeting redundant trending topics, reposted polls, and cross-platform duplicates.
+Subsequently, automated toxicity screening was conducted using the Google Perspective API, with polls retaining a toxicity score below 0.4 being retained to filter out hate speech and overly controversial subjects (Lees et al.
+2022).
   3. 3.
-    Social Relevance Validation: Human verification of voting patterns (removing polls $\leq$ 100 votes for social relevance) and verification of poll and vote existence in all three platforms to ensure authentic social engagement and meaningful community participation.
+    Social Relevance Validation: Human verification of voting patterns (removing polls $\leq$ 100 votes for social relevance) and verification of poll and vote existence in all two platforms to ensure authentic social engagement and meaningful community participation.
   4. 4.
@@
 […]
-We manually annotate each poll includes rich metadata for social context enrichment: general platform context (e.g., user statistics and user behaviors), topical context within that platform (topic-specific discourse patterns), and temporal context for each poll (social events and backgrounds related to the poll when created).
+We manually annotate each poll includes rich metadata for social context enrichment: general platform context (e.g., user statistics and user behaviors) and topical context (topic-specific discourse patterns).
 Cross-Cultural Augmentation.
-We created a parallel bilingual corpus for all three platforms through machine translation with rigorous quality control to demonstrate both linguistic and cultural effects: 10% back-translation validation (BLEU $>$ 35 (papineni2002bleu, )), 5% native speaker rewriting, and expert review (Fleiss’ $\kappa$ = 0.51).
+We created a parallel bilingual corpus for all two platforms through machine translation with rigorous quality control to demonstrate both linguistic and cultural effects: 10% back-translation validation (BLEU >> 35 (Papineni et al.
+2002)), 5% native speaker rewriting, and expert review (Fleiss’ $\kappa$ = 0.51).
 The total number of polls becomes 3,918 after augmentation.
@@
 The final MindVote dataset is composed of 3,918 polls, with 2,158 sourced from Reddit and 1,760 from Weibo.
-Each poll is enriched with a comprehensive set of metadata, including its creation time, total vote count, and three layers of social context: general platform context, topical context within that platform, and the specific temporal context at the time the poll was created.
+Each poll is enriched with a comprehensive set of metadata, including its creation time, total vote count, and layers of social context: general platform context and topical context within that platform.
 To ensure broad accessibility and ease of use, the entire dataset is provided in both CSV and JSON formats.
-Table 1 shows one example with its simplified metadata keywords for illustration.
-Table 1: An example poll from our MindVote dataset, demonstrating the structure of the question, options, and associated social context provided for model evaluation.
+Table 1 shows an example with its simplified metadata keywords.
 Poll: How threatened do you feel by AI replacing your job?
@@
 Topical Context Broad AI adoption. 78% of organizations use AI; 55% of Americans use AI regularly.
-Temporal Context AI-driven layoff spike fuels job insecurity.
+Table 1: An example poll from our MindVote dataset, demonstrating the structure of the question, options, and associated social context provided for model evaluation.
 3.2 Experiment Design
@@
 Models.
-We evaluate 15 leading LLMs across closed-source, open-source and specalization categories: Claude-3.7-Sonnet-02-24 (claude-3.7-thinking, ), GPT-4o-2024-11-20 (openai-4o, ), GPT-4.1-04-14 (openai-4.1, ), Gemini-2.5-Pro-05-06 (google-gemini-2-5-pro, ), o3-medium-04-16 (openai-o3-o4mini-system-card, ), Deepseek-R1-05-28 (2501.12948, ), Qwen2.5-32B-Instruct (qwen2.5, ), Gemma-2-9B-it (team2024gemma, ), Mistral-7B-Instruct-v0.1 chaplot2023albert , Llama-2-13B-Chat touvron2023llama , Llama-3-70B-Instruct (grattafiori2024llama, ), and Llama-4-Maverick-17B-128E (llama4, ).
-We also evaluate three opinion distribution prediction specialization models from suh2025language .
-These models are fine-tuned using LoRA hu2022lora on base models: Llama-3-70B-Base grattafiori2024llama , Llama-2-13B-Base touvron2023llama , and Mistral-7B-v0.1 chaplot2023albert which we named them respectively as: SubPop-Llama-3-70B, SubPop-Llama-2-13B, SubPop-Mistral-7B.
+We evaluate 15 leading LLMs across closed-source, open-source and specialization categories: Claude-3.7-Sonnet-02-24 (Anthropic 2025), GPT-4o-2024-11-20 (OpenAI 2025b), GPT-4.1-04-14 (OpenAI 2025a), Gemini-2.5-Pro-05-06 (Google Cloud 2025), o3-medium-04-16 (OpenAI 2025c), Deepseek-R1-05-28 (DeepSeek-AI et al.
+2025), Qwen2.5-32B-Instruct (Qwen Team 2024), Gemma-2-9B-it (Team, Gemma et al.
+2024), Mistral-7B-Instruct-v0.1 (Chaplot 2023), Llama-2-13B-Chat (Touvron et al.
+2023), Llama-3-70B-Instruct (Grattafiori et al.
+2024), and Llama-4-Maverick-17B-128E (Meta AI 2025).
+We also evaluate three opinion distribution prediction specialization models from Suh et al.
+(2025).
+These models are fine-tuned using LoRA (Hu et al.
+2022) on base models: Llama-3-70B-Base (Grattafiori et al.
+2024), Llama-2-13B-Base (Touvron et al.
+2023), and Mistral-7B-v0.1 (Chaplot 2023) which we named them respectively as: SubPop-Llama-3-70B, SubPop-Llama-2-13B, SubPop-Mistral-7B.
 Pipeline.
-All primary evaluations use a greedy decoding strategy (temperature=0) with default hyperparameter settings under zero-shot with context priming.
-For the specialization fine-tuned models, we adapt those models into our pipeline by first loading the respective pretrained base model and then applying the publicly available LoRA weight checkpoints provided by suh2025language .
+All primary evaluations use a greedy decoding strategy (temperature=0) with default hyperparameter settings under zero-shot with context annotation (Kojima et al.
+2022; Han et al.
+2025).
+For the specialization fine-tuned models, we adapt those models into our pipeline by first loading the respective pretrained base model and then applying the publicly available LoRA weight checkpoints provided by Suh et al.
+(2025).
 Prompt.
 To ensure consistent and machine-readable outputs, we employ a structured prompting strategy where the model is given a JSON object containing the poll and its context.
-The template instructs the model to assume the role of a “opinion distribution prediction expert analyzing voting patterns and social dynamics." The prompt includes the poll question and is enriched with social context metadata served as context priming, with instruction for step-by-step reasoning.
+The template instructs the model to assume the role of a “opinion distribution prediction expert analyzing voting patterns and social dynamics.” The prompt includes the poll question and is enriched with social context metadata, with instruction for step-by-step reasoning.
 The model’s task is to return a JSON object with a schema identical to the input, replacing placeholder fields with its numeric predictions for the voting distribution.
-The complete and structured template is detailed in Appendix LABEL:appendix:prompt.
 Evaluation Metrics.
 We adopt four distinct metrics to provide a comprehensive evaluation.
-Our primary metric is 1 - Wasserstein Distance (1-Wass.) (santurkar2023whose, ; meister-etal-2025-benchmarking, ; suh2025language, ).
+Our primary metric is 1 - Wasserstein Distance (1-Wass.) (Santurkar et al.
+2023; Meister, Guestrin, and Hashimoto 2025; Suh et al.
+2025).
 The Wasserstein Distance measures the minimum cost for transforming one distribution into another, crucially accounting for semantic similarity between answer choices by treating them as points in a metric space.
-To complement this, we also report Spearman’s Rank Correlation Coefficient ($\rho$) (zhou-etal-2022-distributed, ; pavlopoulos2023distance, ), a non-parametric measure of how well the predicted ranking of options matches the true ranking of vote shares; 1 - KL Divergence (meister-etal-2025-benchmarking, ; nguyen2025task, ), which quantifies the information loss when using the model’s predicted distribution to approximate the ground truth; and One-hot Accuracy (zhou-etal-2022-distributed, ; santurkar2023whose, ; suh2025language, ), which provides a strict measure of whether the single most likely predicted answer is correct.
+To complement this, we also report Spearman’s Rank Correlation Coefficient ($\rho$) (Zhou, Nie, and Bansal 2022; Pavlopoulos and Likas 2023), a non-parametric measure of how well the predicted ranking of options matches the true ranking of vote shares; 1 - KL Divergence (Meister, Guestrin, and Hashimoto 2025; Nguyen et al.
+2025), which quantifies the information loss when using the model’s predicted distribution to approximate the ground truth; and One-hot Accuracy (Zhou, Nie, and Bansal 2022; Santurkar et al.
+2023; Suh et al.
+2025), which provides a strict measure of whether the single most likely predicted answer is correct.
 Evaluation Boundary.
-We include upper bounds and lower bounds for comparisons following suh2025language .
+We include upper bounds and lower bounds for comparisons following (Suh et al.
+2025).
 The upper bound is established by sampling subsets of the original results, calculating the four metrics between subsampled and original distributions, and performing bootstrapping to obtain a robust estimate that captures the intrinsic variance arising from the respondent sampling process in opinion.
 The uniform distribution lower bound establishes a performance floor equivalent to random chance.
-Table 2: Opinion distribution prediction performance of LLMs on the MindVote Benchmark.
-Scores are presented as Mean values, evaluated on four different metrics: 1-Wasserstein distance (1-Wass.), 1-KL Divergence (1-KL Div.), Spearman’s Rank Correlation (Spearman.), and One-hot Accuracy (Acc.).
-All metrics are the higher the better.
+4 Results
+4.1 Performance of General Purpose Models
+We analyze the overall performance of ten leading general purpose LLMs on the MindVote benchmark in Table 2.
+Across all four metrics, the closed-source model o3-medium consistently outperforms all other models, establishing the highest performance ceiling among current general-purpose LLMs.
+Among open-source models, Deepseek-R1 demonstrates the strongest results, narrowly outperforming other models in its class.
+However, a critical gap remains between the top-performing models and the upper bound for all metrics.
+This highlights substantial opportunities for improvement in modeling authentic public opinion distribution.
+4.2 Survey-based Specialization Pitfalls
+Our results in Table 2 show that fine-tuning on a structured survey dataset leads to a significant decrease in performance when predicting opinion distributions on social media.
+To demonstrate this, we compare our specialization models (base models that fine-tuned on traditional but context-stripped survey data) against their original base instruction-tuned models (the general-purpose models).
+As shown in Table 2, specialization models consistently underperform their base counterparts, even when evaluated with full context annotations.
+For example, the SubPop-Llama-2-13B specialization model experiences a 3.3 percentage point decrease in its 1-Wass.
+score relative to the base Llama-2-13B.
+These results indicate that fine-tuning on sanitized data does not equip models to handle the complexities of real-world contexts.
+Rather than enhancing performance, such narrow specialization actually impairs the models’ ability to generalize to authentic social discourse.
 Model 1-Wass. 1-KL Div. Spearman. Acc.
@@
 Claude-3.7-Sonnet 0.891 0.851 0.722 0.551
-GPT-4o 0.880 0.836 0.691 0.711
-GPT-4.1 0.874 0.845 0.688 0.697
+GPT-4o 0.880 0.836 0.691 0.515
+GPT-4.1 0.874 0.845 0.688 0.524
 Open-source Models
@@
 Lower Bound 0.701 0.663 0.000 0.307
-4 Result
-4.1 Performance of General Purpose Models
-We analyze the overall performance of ten leading general purpose LLMs on the MindVote benchmark in Table 2.
-Across all four metrics, the closed-source model o3-medium consistently outperforms all other models, establishing the highest performance ceiling among current general-purpose LLMs.
-Among open-source models, Deepseek-R1 demonstrates the strongest results, narrowly outperforming other models in its class.
-However, a critical gap remains between the top-performing models and the upper bound for all metrics.
-This highlights substantial opportunities for improvement in modeling authentic public opinion distribution.
-4.2 Survey-based Specialization Pitfalls
-Our result in Table 2 reveals that fine-tuning on a structured survey dataset results in a significant performance drop on authentic, in-domain tasks.
-To demonstrate this, we compare our specialization models (base models that fine-tuned on traditional but context-stripped survey data) against their original base instruction-tuned models (the general-purpose models).
-As shown in Table 2, the specialization models consistently underperform their base counterparts.
-This evaluation was conducted under our standard testing condition, where all models were provided with full context served as context priming setting.
-For instance, the SubPop-Llama-2-13B specialization model exhibits a 3.3 percentage point drop in its 1-Wass.
-score compared to the base Llama-2-13B model.
-This finding suggests that fine-tuning on sanitized data fails to equip models with the necessary capabilities to handle real-world contextual complexity.
-Instead of improving performance, this narrow specialization is detrimental, degrading the models’ ability to generalize to authentic social discourse.
+Table 2: Opinion distribution prediction performance of LLMs on the MindVote Benchmark.
+Scores are presented as Mean values, evaluated on four different metrics: 1-Wasserstein distance (1-Wass.), 1-KL Divergence (1-KL Div.), Spearman’s Rank Correlation (Spearman.), and One-hot Accuracy (Acc.).
+All metrics are the higher the better.
 4.3 Performance Across Topics and Cultures
 Performance varies by Topic.
-As shown in Table 3, model performance is not uniform across different topical domains.
-On average, models achieve higher performance on Social Issues and Lifestyle topics, which often feature more structured discourse patterns that align with standard pre-training corpora.
-In contrast, performance tends to be lower in domains like Technology and Entertainment where specialized jargon, community-specific norms, and rapidly evolving vernacular are more common.
-This disparity suggests that models struggle to adapt their reasoning from general knowledge to the specific linguistic and social conventions of diverse online communities.
-Performance Reflects Cultural Origin.
-The analysis in Figure 4 reveals a ’Cultural Gap,’ showing models have a strong cultural alignment with their origin.
-Western models like Gemini-2.5-Pro excel on Reddit, while Chinese models such as DeepSeek-R1 dominate on Weibo.
-We confirmed this gap is primarily cultural, not linguistic, by evaluating translated content; the performance penalty from translation was minimal compared to the large drop across cultural sources.
-This systematic variation demonstrates a clear alignment effect, where models are more adept at interpreting the cultural norms of their training data.
-4.4 Analysis
-Section 4.2 reveals a critical survey-based specialization pitfall: base models fine-tuned on sanitized survey datasets exhibit degraded performance even with context priming.
-This unexpected finding raises three questions that we expect to investigate:
-  • (1) What role does social context actually play?
-  • (2) Why do models struggle with contextual complexity?
-  • (3) Is it more beneficial for a model to understand specific context or to imitate examples?
-These questions emerge naturally from our specialization findings.
-The failure of fine-tuned models despite having full context suggests either that platform, topic and temporal context are less important than assumed, or that these models cannot effectively utilize those contextual information.
-This motivates our first investigation into whether social context is truly beneficial signal or noise.
-The stark performance gap between base and specialized models further suggests that certain contextual characteristics may be particularly challenging, leading to our second analysis of contextual complexity dimensions.
-Finally, the contrast between our context-rich approach and traditional few-shot learning methods—where specialized models were trained on examples rather than contextual understanding—compels us to directly compare these two information-providing strategies.
-Table 3: Opinion distribution prediction performance across different topics. The final row shows the average performance across all models for each major topic.
+As shown in Table 3, model performance differs notably across topical domains.
+On average, models perform better on Social Issues and Lifestyle topics, which typically exhibit more structured discourse aligned with standard pre-training corpora.
+In contrast, performance declines in domains such as Technology and Entertainment, where specialized jargon, community-specific norms, and rapidly evolving vernacular are prevalent.
+This disparity indicates that models have difficulty adapting their reasoning from general knowledge to the nuanced linguistic and social norms of diverse online communities.
 Model Tech. Soc. Enter. Career Life.
@@
 Average 0.830 0.863 0.838 0.844 0.857
-Social Context is a Critical Signal.
-Our context ablation study (Table 4) confirms that social context is not noise but a vital signal for accurate opinion prediction.
-Removing all contextual metadata (No Context) results in the most substantial performance drop on average (-5.91%), followed closely by platform-specific information removal (-5.12%) and topical context removal (-4.52%).
-Temporal context shows the smallest impact, but its existence is important to provide backgrounds for questions need temporal information.
-Notably, specialized fine-tuned models exhibit dramatically larger degradation across all ablation conditions.
-This severe degradation approaches the theoretical lower bound of uniform distribution performance, suggesting these models have become overly dependent on contextual signals.
-Since these models were fine-tuned on structured survey data in standardized formats, they struggle to adapt when deployed on natural social media polls when contextual information may be fragmented or entirely absent.
-While these models may excel in controlled survey environments, their heightened sensitivity to context removal in real-world social media settings reveals a limitation, potentially constraining their applicability across diverse online environments.
+Table 3: Opinion distribution prediction performance across different topics. The final row shows the average performance across all models for each major topic.
+Performance Reflects Cultural Origin.
+The analysis in Figure 4 reveals a Cultural Gap, showing models have a strong cultural alignment with their origin.
+Western models like Gemini-2.5-Pro excel on Reddit, while Chinese models such as DeepSeek-R1 dominate on Weibo.
+We confirmed this gap is primarily cultural, not linguistic, by evaluating translated content; the performance penalty from translation was minimal compared to the large drop across cultural sources.
+This systematic variation demonstrates a clear alignment effect, where models are more adept at interpreting the cultural norms of their training data.
+4.4 Analysis
+Section 4.2 highlights a pitfall of survey-based specialization: base models fine-tuned on sanitized survey datasets show degraded performance, even when context priming—adding relevant situational or background information to guide predictions—is applied (Doyle and Lee 2016).
+This unexpected finding raises three questions that we expect to investigate:
+  • Does social context in social media serve as noise or a helpful signal?
+  • How does increasing contextual complexity affect model performance, especially for survey-specialized models?
+  • Does adding social contexts better enhance models’ opinion prediction compared with few-shot learning?
+These questions emerge naturally from our specialization findings.
+The failure of fine-tuned models, even when provided with full context, suggests two possibilities: either (1) platform and topic are less important than previously assumed, or (2) these models are unable to effectively utilize such contextual information.
+This motivates our first investigation into whether social context is truly beneficial signal or noise.
+Moreover, the suboptimal performance of specialized models reflects their difficulty in processing social context, motivating our second analysis on factors contributing to context complexity.
+Third, because our context-rich approach relies on contextual understanding rather than example-based training as in few-shot learning, a direct comparison is needed to assess the effectiveness of both strategies.
 Figure 4: LLMs exhibit a cultural alignment with their origin.
@@
 Error bars represent 95% CI.
-Contextual Complexity.
-Our ablation study showed that removing social context severely degrades performance.
-To understand why this context is so critical, we investigate how its inherent complexity impacts model behavior.
-We define contextual complexity using three metrics: (1) character length of combined metadata; (2) language informality (measured by slang and emoji frequency); and (3) topic niche nature (inversely proportional to dataset percentage).
-As demonstrated in Figure 5, our findings reveal consistent negative correlations between model performance and all complexity metrics across evaluated systems.
-This indicates that as social contexts become longer, more informal, and more specialized, models systematically exhibit reduced prediction certainty.
-The analysis reveals distinct performance tiers, with closed-source models showing greater resilience, open-source models displaying moderate sensitivity, and specialization models exhibiting the most pronounced degradation under complexity.
-This pattern suggests that domain-specific fine-tuning on formal survey data creates brittleness when encountering the informal, less structured nature of social media polls.
-The hierarchical performance differences confirm that authentic social discourse’s rich, messy characteristics pose universal challenges for LLMs, with models fine-tuned on structured formats being particularly vulnerable to contextual complexity.
-Table 4: Opinion distiribution prediction performance degradation from the full-context baseline.
+Social Context is a Critical Signal.
+Our context ablation study (Table 4) confirms that social context is not noise but a vital signal for accurate opinion prediction.
+Removing all contextual metadata (No Context) results in the most substantial performance drop on average (-5.91%), followed closely by platform-specific information removal (-5.12%) and topical context removal (-4.52%).
+Notably, specialized fine-tuned models exhibit dramatically larger degradation across all ablation conditions.
+This severe degradation approaches the theoretical lower bound of uniform distribution performance, suggesting these models have become overly dependent on contextual signals used for priming.
+Since these models were fine-tuned on structured survey data in standardized formats, they struggle to adapt when deployed on natural social media polls when contextual information may be fragmented or entirely absent.
+While these models may excel in controlled survey environments, their heightened sensitivity to context removal in real-world settings reveals a limitation, constraining their applicability across diverse online environments.
+Model w/o Plat. w/o Topic. No Ctx.
+Closed-source Models
+o3-medium -3.03 -2.31 -4.13
+Gemini-2.5-Pro -2.37 -3.30 -4.34
+Claude-3.7-Sonnet -4.46 -3.91 -5.53
+GPT-4o -5.75 -3.39 -4.92
+GPT-4.1 -4.92 -3.36 -5.30
+Open-source Models
+Deepseek-R1 -4.19 -3.55 -5.98
+Qwen2.5-32B -4.42 -3.81 -6.24
+Llama-3-70B -4.82 -4.78 -5.19
+Llama-4-17B -5.80 -5.33 -6.08
+Llama-2-13B -4.95 -5.26 -5.92
+Gemma-2-9B -5.37 -4.74 -6.47
+Mistral-7B -5.92 -4.84 -6.55
+Specialization Models
+SubPop-Llama-3-70B -6.82 -6.68 -6.95
+SubPop-Llama-2-13B -6.94 -6.11 -7.23
+SubPop-Mistral-7B -6.89 -6.41 -7.54
+Average -5.12 -4.52 -5.91
+Table 4: Opinion distribution prediction performance degradation from the full-context baseline.
 All scores represent the drop in 1-Wasserstein Distance (%).
@@
 (Platform), Topic.
-(topical), Temp.
-(Temporal), Ctx.
+(topical), Ctx.
 (Context).
-Model w/o Plat. w/o Topic. w/o Temp. No Ctx.
-Closed-source Models
-o3-medium -3.03 -2.31 -1.02 -4.13
-Gemini-2.5-Pro -2.37 -3.30 -0.98 -4.34
-Claude-3.7-Sonnet -4.46 -3.91 -1.39 -5.53
-GPT-4o -5.75 -3.39 -0.87 -4.92
-GPT-4.1 -4.92 -3.36 -1.57 -5.30
-Open-source Models
-Deepseek-R1 -4.19 -3.55 -1.04 -5.98
-Qwen2.5-32B -4.42 -3.81 -0.88 -6.24
-Llama-3-70B -4.82 -4.78 -2.11 -5.19
-Llama-4-17B -5.80 -5.33 -1.01 -6.08
-Llama-2-13B -4.95 -5.26 -1.37 -5.92
-Gemma-2-9B -5.37 -4.74 -1.26 -6.47
-Mistral-7B -5.92 -4.84 -1.46 -6.55
-Specialization Models
-SubPop-Llama-3-70B -6.82 -6.68 -1.88 -6.95
-SubPop-Llama-2-13B -6.94 -6.11 -2.10 -7.23
-SubPop-Mistral-7B -6.89 -6.41 -2.96 -7.54
-Average -5.12 -4.52 -1.46 -5.91
+Contextual Complexity.
+While our ablation study confirms that social context is critical for performance, its inherent complexity presents a significant challenge.
+To dissect this, we investigate how model performance degrades as complexity increases about our annotated context.
+We define contextual complexity using three metrics: (1) context length by tokens of combined metadata; (2) language informality in the context; and (3) niche topic which constitutes a small fraction of the overall dataset.
+Our analysis in Figure 5 reveals that increasing complexity in any of these areas has negative correlations with the 1-Wass.
+performance.
+This effect is most pronounced for our specialization models.
+Specifically, these models, fine-tuned on structured and formal survey data, exhibit strong performance degradation when faced with longer contexts, higher degrees of informality, and more niche topics.
+This demonstrates a critical brittleness: while trained to be domain experts, their reliance on formal data structures makes them particularly vulnerable to the unstructured, multifaceted, and messy nature of authentic social media discourse.
+Figure 5: Correlation patterns between model’s 1-Wass. performance and complexity dimensions. Survey-specialized models exhibit strong brittleness to social media contextual complexity.
 Contextual Priming vs.
 Few-Shot Learning.
-Our “survey-based specialization pitfall" finding (Section 4.2) raised a key question: do models benefit more from understanding a poll’s environment via contextual priming or from imitating examples via few-shot learning?
-To test this, we compared model performance in a zero-shot setting with full context against a zero-shot setting with no context and a few-shot setting with no context.
-The results in Figure 6 are stark: contextual priming yields a substantial increase across all models, exceeding the impact of few-shot examples in most cases.
-Moreover, few-shot performance shows instability across models but contextual priming provides more consistent improvements.
-This finding suggests that the ability to situate a problem within its social context is a more critical capability than pattern-matching from isolated examples.
-Figure 5: Performance degradation patterns across model types and complexity dimensions.
-The heatmap reveals a clear hierarchy: closed-source models maintain robustness while survey-specialized models exhibit severe brittleness to social media contextual complexity.
+Our “survey-based specialization pitfall” finding (Section 4.2) raised a key question: do models benefit more from understanding a poll’s environment via contextual priming or from imitating examples via few-shot learning?
+To evaluate this, we compared the model’s performance across three distinct settings.
+The first was a zero-shot setting with context priming, which followed our default context annotation pipeline.
+We contrasted this against two settings without context priming: a standard zero-shot setting without context priming, same as we have done in Table 4 and a few-shot (1-4 examples) setting.
+For the few-shot configuration, we supplied the model with examples predicted from Claude-3.7, where each example contained a poll, its distribution prediction, and the corresponding reasoning steps.
+The results in Figure 6 are stark: contextual priming delivers a substantial performance boost across all models, often exceeding the gains achieved through few-shot learning.
+In contrast, the benefit from few-shot examples is unreliable.
+As the figure shows, performance does not reliably increase with the number of examples; for instance, the average 4-shot performance of open-sourced models is worse than that of 3-shot performance.
+This discrepancy suggests that the ability to situate a problem within its broader social context is a more critical and robust capability than simple pattern-matching from isolated examples.
+Figure 6: Contextual priming outperforms few-shot learning across all model categories, highlighting the importance of social context over example-based pattern matching.
 5 Error Analysis
 We conduct an error analysis to categorize prediction reasoning errors  [^1] We consider the individual poll’s 1-Wass.
-$<$ 0.8 as error..
+<< 0.8 as error..
 Using the LLM-as-a-judge, we identify three main categories of error type shown in Table 5.
-Figure 6: Contextual priming outperforms few-shot learning across all model categories, highlighting the importance of social context over example-based pattern matching.
-Table 5: Analysis of Claude 3.7 Sonnet prediction errors categorized by failure modes: Platform Mis-Adaptation (42.5%) – misapplying knowledge about specific platforms; Cultural Misalignment (36.6%) – ignoring local and cultural contexts; Temporal Dislocation (20.0%) – mis-understanding about the temporal context.
 Error Type Case Study & Failed Reasoning
@@
 Cultural Misalignment New phone Buy or Wait (Weibo): Assumed universal tech-enthusiasm drives upgrades, overlooking Chinese market’s specific socio-economic factors and consumer behavior patterns.
-Temporal Dislocation Calling it "X" vs. "Twitter" (Reddit): Reasoning anchored to official rebrand timeline, underestimating persistent colloquial usage and real-world adoption resistance.
+Temporal Dislocation Calling it ”X” vs. ”Twitter” (Reddit): Reasoning anchored to official rebrand timeline, underestimating persistent colloquial usage and real-world adoption resistance.
+Table 5: Analysis of Claude 3.7 Sonnet prediction errors categorized by failure modes: Platform Mis-Adaptation (42.5%) – misapplying knowledge about specific platforms; Cultural Misalignment (36.6%) – ignoring local and cultural contexts; Temporal Dislocation (20.0%) – mis-understanding about the temporal information.
 Platform Mis-Adaptation.
-A model can fail into platform mis-adaptation, where they misapplying their knowledge about certain platform, such as when analyzing a Reddit poll about paid TikTok subscriptions.
-It may apply a generic economic lens, predicting high tendency toward worth option based on the influencer economy.
+In the TikTok subscriptions case, the model applies a generic economic lens, predicting high tendency toward worth option based on the influencer economy.
 This reasoning ignores the context of Reddit, where the user tend to be negative of influencer monetization.
@@
 Cultural Misunderstanding.
-In the phone replacement case, models systematically misinterpret Chinese users’ pragmatic spending attitudes, failing to capture the genuine cost-conscious preferences common among Weibo users, which often contradict Western-centric consumer marketing narratives.
+In the phone update case, models systematically misinterpret Chinese users’ pragmatic spending attitudes, failing to capture the genuine cost-conscious preferences common among Weibo users, which often contradict Western-centric consumer marketing narratives.
 While model recognizes about the technical enthusiasm in Weibo, their predictions fail to reflect the cost-conscious approach that characterizes technology adoption for many Chinese social media users.
 Temporal Dislocation.
-This error happens when models fail to account for the lag between official changes and actual user adoption patterns.
 In the Twitter-to-X rebranding case, the model assumes that users would quickly adopt the new term.
@@
 6 Conclusion
-We introduce MindVote, the first benchmark for public opinion prediction grounded in authentic social media discourse.
+We introduce MindVote, the first benchmark for public opinion distribution prediction in social media.
 Our comprehensive evaluation demonstrates the critical importance of assessing models in naturalistic, context-rich environments.
@@
 MindVote provides the essential tool to guide and measure this necessary shift.
+7 Limitation
+While MindVote represents a significant advance in ecological validity, our benchmark is limited to two platforms, Reddit and Weibo, which, despite their scale and cultural diversity, do not capture the full spectrum of social media ecosystems.
+Platforms such as Twitter/X, TikTok, Instagram, and region-specific networks exhibit distinct interaction patterns, demographic compositions, and content modalities that may influence opinion formation differently.
+Future work should expand to additional platforms to ensure comprehensive evaluation of LLMs’ social reasoning capabilities across diverse digital environments.